OpenStudy (anonymous):

a bag contains 4 green marbles and 5 white marbles. suppose a marble is randomly selected. what are the odds in fvor of picking a white marble?

OpenStudy (anonymous):

the ratio of the number of white marbles to the number of not white marbles

OpenStudy (anonymous):

odds in favor is p(event) divided by 1-p(e)

OpenStudy (anonymous):

9 marbles. probability of drawing a white is 5/9

OpenStudy (anonymous):

i guess in english it would be the ratio of the white marbles to the green marbles

OpenStudy (anonymous):

so \[\frac{\frac{5}{9}}{1-\frac{5}{9}}\]

OpenStudy (anonymous):

i.e. "number of white marbles : number of green marbles"

OpenStudy (anonymous):

you can do it @mathbrz way too i prefer not to mess with the fractions just count

OpenStudy (anonymous):

yeah simple and effective is better lol